How many pounds of coffee worth $1.80 a pound must be mixed with 15 pounds of coffee worth $1.10 a pound to produce a mixure worth $1.50 a pound?
Answer by scott8148(6628)

x(1.80) + 15(1.10) = (x+15)(1.50)

1.8x + 16.5 = 1.5x + 22.5

.3x = 6 ___ x = 20